OpenStudy (anonymous):

√108 - 3√50 + 5√12 + 4√18 I'm kind of confused on how to simplify..

jimthompson5910 (jim_thompson5910):

Notice how \[\Large \sqrt{108} = \sqrt{36*3}\] \[\Large \sqrt{108} = \sqrt{36}*\sqrt{3}\] \[\Large \sqrt{108} = 6\sqrt{3}\]

jimthompson5910 (jim_thompson5910):

I'm factoring the number into two numbers where one number is a perfect square

jimthompson5910 (jim_thompson5910):

Then I'm breaking up the root and simplifying

jimthompson5910 (jim_thompson5910):

Use this idea to simplify the other roots

OpenStudy (anonymous):

Would the second one be,

OpenStudy (anonymous):

3√(25*2)=3√25*√2=3*5√2

jimthompson5910 (jim_thompson5910):

which becomes \[\Large -15\sqrt{2}\]

OpenStudy (anonymous):

Why -15?

jimthompson5910 (jim_thompson5910):

because it was -3 instead of +3

jimthompson5910 (jim_thompson5910):

in front of the square root of 50

OpenStudy (anonymous):

What about the last one.. 4√18

jimthompson5910 (jim_thompson5910):

hold on, what did you get for 5√12

jimthompson5910 (jim_thompson5910):

did you manage to simplify that at all?

OpenStudy (anonymous):

For 5√12, I got 10√3 ?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

for 4√18 factor 18 into 9*2, then simplify

OpenStudy (anonymous):

Thank you!

jimthompson5910 (jim_thompson5910):

you're welcome, tell me what your final answer is

OpenStudy (anonymous):

6√3 - 5√2 + 10√3 + 12√2

OpenStudy (anonymous):

i mean, -15√2

jimthompson5910 (jim_thompson5910):

√108 - 3√50 + 5√12 + 4√18 should be 6√3 - 15√2 + 10√3 + 12√2 Notice it's -15√2 and not -5√2

jimthompson5910 (jim_thompson5910):

Now combine like terms

OpenStudy (anonymous):

16√3 - 3√2

jimthompson5910 (jim_thompson5910):

good, you nailed it

OpenStudy (anonymous):

[:

jimthompson5910 (jim_thompson5910):

Another way to verify is to type in (√108 - 3√50 + 5√12 + 4√18) - (16√3 - 3√2) into your calculator If you get the result of 0 (or very very close to it), then you have verified the answer